OpenStudy (anonymous):

Explain: log(x+3/10)+log x+1 =0

OpenStudy (experimentx):

log(x+3/10)+log(x+1) =0 ??

OpenStudy (anonymous):

^^yes

OpenStudy (anonymous):

explain: use the log rule so you can rewrite the left side as the log of a quotient. since the bases are the same for the logs, the arguments are the same... solve...

OpenStudy (experimentx):

log((x+3)/(x+1))

OpenStudy (anonymous):

u good? i'm good... :)

OpenStudy (experimentx):

looks like a typo there ** multiplication

OpenStudy (experimentx):

log(x+3) - log(10) + log(x+1) = 0 log((x+3)*(x+1)) = log(10) (x+3)*(x+1) = 10

OpenStudy (anonymous):

yep. @experimentX , is correct.... how did i miss that?

OpenStudy (anonymous):

i'm not good

OpenStudy (experimentx):

what have you been smokin buddy ... thats not like you!!

OpenStudy (anonymous):

i've been at level 0 for several times today... :(

OpenStudy (experimentx):

must be a bug.

OpenStudy (anonymous):

log(x+3) - log(10) + log(x+1) = 0 log((x+3)*(x+1)) = log(10) (x+3)*(x+1) = 10 x^2 + 4x + 3 = 10 x^2 + 4x - 7 = 0 quadratic formula...

OpenStudy (experimentx):

some maniac gave me ... i don't know how much medals ... i was at level 60 when i woke up i was at level 62. I messaged "why are you overdoing it?" I got reply ""[Hitler] has grasped the falsity of the hedonistic attitu ..... a result a whole nation flings itself at his feet." ...... -(George Orwell's review of Mein Kampf) Hopefully that will suffice."

OpenStudy (experimentx):

a page long essay .. lol.

OpenStudy (anonymous):

wow...

OpenStudy (anonymous):

I am still confused at to which answer I should go with here.. (Forgive my beginnerness)

OpenStudy (anonymous):

sorry... you'll need to solve that last quadratic equation: x^2 + 4x - 7 = 0. Since this is not factorable, you can use the quadratic formula.

OpenStudy (anonymous):